OpenStudy (anonymous):

REALLY NEED HELP ALGEBRA 1!!!! CONNNECTIONS?? Using the discriminant, how many real number solutions does this equation have?

OpenStudy (perl):

whats the equation

OpenStudy (anonymous):

3x2 – 2 = 5x

OpenStudy (anonymous):

@perl sorry lol

OpenStudy (is3535):

prel i need help

OpenStudy (perl):

$$ \Large \rm{ ax^2 + bx + c = 0 \\~\\ \begin{cases} b^2 -4ac > 0~~two~ real ~root ~ \\ b^2 -4ac =0~~one~ real ~root ~ \\ b^2 -4ac <0 ~~ two ~non ~real \end{cases} }$$

OpenStudy (anonymous):

so is there 2?

OpenStudy (is3535):

@perl can you help me

OpenStudy (perl):

sorry im helping spanked

OpenStudy (perl):

\[\large {3x^2 – 2 = 5x \\ 3x^2 - 5x - 2 = 0 \\ \rm Discriminant = (-5)^2 - 4(3)(-2) } \]

OpenStudy (perl):

since its positive, you have two real solutions

OpenStudy (anonymous):

Thank you sir! I got it now!

OpenStudy (anonymous):

@perl

OpenStudy (is3535):

@perl